OpenStudy (anonymous):

Factor 6(x^2 - 4x + 4)^2 + (x^2 - 4x + 4) - 1

OpenStudy (anonymous):

um, ok!

OpenStudy (anonymous):

To do this one we have to use the holy crap its hard~!

OpenStudy (anonymous):

I'm listening What is the method called?

OpenStudy (anonymous):

multiplicative property. So 6x^4,- 24x,+ 24, + -1x^2

OpenStudy (anonymous):

um, thats as far as Ive done, Ill finish the problem, however I do not know what to do with the squares.

OpenStudy (anonymous):

I'll probably find out once I finish. The method Im using right now is called the multiplicative property.

OpenStudy (anonymous):

basically 6(x^2 - 4x + 4)^2 +, see this part of the problem? everything in parenthesis needs to be multiplied by 6 individually.

OpenStudy (anonymous):

but before you do that, you need to square that little thing in oh crap... made a bobo..

OpenStudy (anonymous):

You see, I think your suppose to solve the problem in the parenthesis first (obviously) and then square it, and then multiply by 6!

OpenStudy (anonymous):

Hence, without knowing X we will wrtie as ?x^2

OpenStudy (anonymous):

actually, sorry, made a mistake again, you can square x, you just multiply it by itself, if we have 4x nad then we square it, it would give us 16x

OpenStudy (anonymous):

Ok, NOW! having all that cleared up, we can commence finding the factor, or number that goes into this equation... Am I right? I'm a little confused to XD please help. any how, I'll go and try it.

OpenStudy (anonymous):

16 + 4x (6), 16 * 6= 63, 24x, 34x + (x^2 - 4x + 4) - 1 2x-4x=2x+4 (-1) = -2x + -4x 63 + 24x + -2x + -4x

OpenStudy (anonymous):

Ok! were getting really close!

OpenStudy (anonymous):

63 + 24x + -2x + -4x-= 22x + 59

OpenStudy (anonymous):

soooo.....now were stuck with 22x + 58. Is this what you wanted? I'm sorry

OpenStudy (anonymous):

Nope lol, I wanted it broken down into 2 factors.

OpenStudy (anonymous):

WTFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFFF

OpenStudy (anonymous):

oh well, I tried!

OpenStudy (anonymous):

once some one helps you, ill check >.<

OpenStudy (anonymous):

Thank you for trying. I appreciate it.

OpenStudy (anonymous):

welcome :s